# ProbRes: Probabilistic Jump Diffusion for Open-World Egocentric Activity Recognition

### v1

@@ Abstract
Open-world egocentric activity recognition poses a fundamental challenge due to its unconstrained nature, requiring models to infer unseen activities from an expansive, partially observed search space. We introduce ProbRes, a Probabilistic Residual search framework based on jump-diffusion that efficiently navigates this space by balancing prior-guided exploration with likelihood-driven exploitation. Our approach integrates structured commonsense priors to construct a semantically coherent search space, adaptively refines predictions using Vision-Language Models (VLMs) and employs a stochastic search mechanism to locate high-likelihood activity labels while minimizing exhaustive enumeration efficiently. We systematically evaluate ProbRes across multiple openness levels (L0–L3), demonstrating its adaptability to increasing search space complexity. In addition to achieving state-of-the-art performance on benchmark datasets (GTEA Gaze, GTEA Gaze+, EPIC-Kitchens, and Charades-Ego), we establish a clear taxonomy for open-world recognition, delineating the challenges and methodological advancements necessary for egocentric activity understanding. Our results highlight the importance of structured search strategies, paving the way for scalable and efficient open-world activity recognition.
@@ 1 Introduction
Egocentric activity recognition in open-world settings presents a fundamental challenge in perception-driven reasoning, where an intelligent system must infer ongoing activities from a vast, unconstrained space of possibilities. Unlike closed-set classification, where predefined categories limit ambiguity, open-world recognition requires models to adaptively infer plausible activities, even when faced with unknown or unseen combinations of actions and objects. This challenge is particularly relevant for assistive robotics, wearable AI, and autonomous agents, where real-time understanding of human actions is crucial for interaction and decision-making. While multimodal foundation models such as Vision-Language Models (VLMs) have demonstrated remarkable zero-shot generalization, their reliance on exhaustive enumeration for inference makes them inefficient for large-scale open-world reasoning. To bridge this gap, structured knowledge — commonsense priors, search heuristics, and adaptive reasoning — is essential for guiding recognition beyond naive brute-force enumeration, enabling scalable and efficient activity understanding in real-world applications.
@@
The increasing focus on open-world learning [26, 2, 13, 20] has led to ambiguity in defining openness, as different factors—objects, actions, and domains—can independently contribute to it. Without a clear distinction, comparing methods and assessing generalization remain challenging. To address this, we define a hierarchy of openness levels that systematically capture increasing uncertainty in the search space (Figure 1). At L0, as in traditional zero-shot learning, all possible activity categories (action-object combinations) are predefined, but some may lack training examples. This setting benefits directly from Vision-Language Models (VLMs), which can generalize to unseen instances but operate within a bounded conceptual space. At L1, the atomic concept space — objects and actions — is known, but their composition (activities) is undefined, requiring models to reason about plausible pairings. At L2, the domain is known (e.g., cooking), but the underlying search space is unconstrained. Finally, at L3, no prior knowledge about the domain, actions, or objects is available, requiring models to construct the entire search space on the fly. As openness increases, exhaustive enumeration becomes intractable, necessitating structured priors and adaptive search strategies.
@@
To address these, we introduce Probabilistic Residual Search (ProbRes), a structured search framework designed to navigate unconstrained activity spaces efficiently. Instead of relying solely on VLM-based likelihood estimation, ProbRes integrates structured priors from external knowledge sources to guide exploration, ensuring semantically meaningful search trajectories without supervision. As illustrated in Figure 2, ProbRes operates in three key phases: (1) exploration, where prior-driven sampling identifies plausible activity candidates; (2) exploitation, where the search refines predictions based on VLM likelihoods; and (3) residual refinement, which further optimizes results by re-ranking activities based on their decomposed action-object components. By balancing structured priors with data-driven reasoning, ProbRes efficiently locates high-likelihood activities while drastically reducing VLM queries for scalable open-world egocentric activity recognition.
@@
Our contributions are four-fold: (i) we propose ProbRes, a probabilistic residual search strategy that integrates structured priors with likelihood-driven inference, effectively balancing exploration and exploitation for open-world activity recognition, (ii) we introduce a clear hierarchical taxonomy (L0–L3) to systematically define and evaluate open-world recognition challenges, providing a principled foundation for future research, (iii) we show that ProbRes drastically reduces VLM queries while maintaining or improving recognition accuracy, demonstrating the importance of structured priors in large-scale search, and (iv) we highlight the limitations of current VLM-based text embeddings for search efficiency with qualitative analysis and highlight the need for improved semantic structuring in open-world reasoning.
@@ 2 Related Work
Egocentric video understanding has gained significant traction in the computer vision community in recent years due to its critical applications in areas such as imitation learning for robotic skills [45], virtual reality [21], and autonomous skill acquisition [22, 27, 4]. A wide range of tasks have been proposed to analyze egocentric video understanding, including procedure understanding, EgoPose [18, 19], proficiency estimation, question answering [15], gaze prediction [1, 29, 16], activity recognition [28], and summarization [33]. The development of large-scale datasets such as EgoExo4D [19], Ego4D [18], EPIC-Kitchens [10], Charades-Ego [46], GTEA Gaze [29], GTEA Gaze Plus [16], and Assembly101 [44] has been pivotal to the progress.
@@
Despite these advancements, action recognition in egocentric videos remains challenging, particularly under conditions of rapid camera movement and significant occlusion, leading to issues with accuracy and generalization across different environments. While early approaches to egocentric action recognition were dominated by supervised learning methods—such as spatial-temporal modeling [49], motion-based representations [34], hand-object interactions [53, 62, 56], and time-series modeling [43]—recent research has shifted towards self-supervised learning [59, 57], leveraging large-scale pretraining followed by fine-tuning for downstream tasks. Both paradigms remain data-driven and require extensive training to achieve optimal performance. Multi-modal LLM-based models such as MM-Ego [58] and GPT4Ego [9] have recently been explored for egocentric action recognition. Zero-shot learning [46, 60] offers a potential avenue to reduce dependence on labeled datasets, but its effectiveness is constrained by the inherent limitations of pre-trained models in reasoning about open-world activity compositions.
@@
Open-world understanding, particularly in egocentric activity recognition, remains an underexplored area. While significant progress has been made in open-world object detection, such as Open World DETR [13] and open-vocabulary object detection [20, 14], which leverage VLMs through prompting and knowledge distillation, challenges remain in handling novel class biases, as highlighted by UMB [55]. One of the first works addressing open-world egocentric activity recognition, KGL [2], employed a neuro-symbolic framework utilizing ConceptNet [47] for knowledge-driven inference. However, its reliance on object detectors for grounding concepts and the semantic gap between knowledge bases and video data limits its applicability. Chatterjee et al. [7] proposed a verb encoder and prompt-based object encoder for open-vocabulary action recognition in egocentric videos. More recently, ALGO [26] extended neuro-symbolic prompting to refine action-object affinities iteratively, combining object grounding with video foundation models. Other works explore related directions, such as incremental open-world action recognition [39] and skeleton-based methods for open-set action recognition [36]. Despite these advances, existing approaches struggle in unconstrained search spaces.
@@
Vision-Language Models have gained prominence following the success of transformer-based LLMs such as BERT [12], RoBERTa [32], GPT series [40, 41, 5, 6], and open-source models like LLaMA [51, 52, 17] and DeepSeek [11]. Image-language models, including CLIP [42], DeCLIP [30], and ALIGN [23], leverage large-scale image-text pairs with contrastive learning [8, 24] to achieve strong zero-shot classification. Video-language models extend these capabilities to egocentric video understanding. EGO-VLP [31] employs separate video and text encoders, Hier-VL [3] captures multi-scale video information hierarchically and LAVILA [61] enhances video-text embeddings with LLM-generated narrations. EgoVLPv2 [38] integrates cross-modal fusion, while ALANAVLM [50] advances large-scale multimodal pretraining for embodied AI. However, these models are inherently constrained by the fixed search space defined during training, limiting their generalization ability in open-world scenarios. In this work, we introduce ProbRes, which dynamically explores and refines activity predictions beyond predefined label sets by integrating structured priors and adaptive search mechanisms.
@@ 3 ProbRes: Jump Diffusion-based Reasoning
Overview. We aim to infer an activity label $a\in\mathcal{S}$ from a given video observation $v$. Unlike conventional closed-set recognition [49], where all activities are known during training, open-world settings require reasoning over a vast, partially observed search space, making exhaustive enumeration computationally prohibitive. Given a prior probability distribution $P_{\text{prior}}(a)$ capturing activity occurrence, a Vision-Language Model (VLM) likelihood estimator $P_{\text{likelihood}}(v{\mid}a)$, and a search space $\mathcal{S}$, the goal is to estimate the most probable activity efficiently:
@@
5:         $a_{t+1}\sim P_{\text{explore}}(a)$triangleright Sample with exploration
6:     $\triangleright$ Exploitation Phase
@@
8:         $a_{t+1}\sim P_{\text{guided}}(a)$triangleright Sample with exploitation
@@
12:$\mathcal{A}_{\text{refine}}=\arg\max_{a\in\mathcal{S}}P_{\text{likelihood}}(v\mid a)$triangleright Top candidates for refinement
@@
14:     $a\rightarrow(a_{\text{action}},a_{\text{object}})$triangleright Decompose activity
15:     $S_{a}=v^{T}\phi(a_{\text{action}})$triangleright Action Likelihood
16:     $S_{o}=v^{T}\phi(a_{\text{object}})$triangleright Object Likelihood
@@ 3.1.2 Search Space Structuring
The search is performed in the text embedding space of activities, where each candidate is represented by a Vision-Language Model (VLM) embedding $\phi(a)$. However, raw VLM embeddings do not always encode the fine-grained affordance relationships necessary for effective search. We construct an ordered search space to ensure that the search moves in semantically meaningful directions. First, we extract action-object phrase embeddings from the VLM to form an initial unordered set $\mathcal{S}=\{a_{i}=(a_{\text{action}},a_{\text{object}})\}$. We define a similarity metric based on the Euclidean distance between embeddings, $d(a_{i},a_{j})=||\phi(a_{i})-\phi(a_{j})||$, where $\phi(a)$ represents the embedding of activity $a$. The search space is structured by selecting an anchor point $a_{\text{ref}}$ and sorting activities based on their distance to $a_{\text{ref}}$. This ensures that similar activities remain proximate, enabling ProbRes to perform jumps that respect semantic locality.
@@ 3.2.3 Concept Decomposition and Re-Ranking
Implementation Details. We use EGOVLP [31] and LAVILA [61] as VLM backbones for likelihood estimation, guided by structured priors from ConceptNet. For L2 and L3 evaluations, search spaces were generated using Gemini 2.0 Flash [37] and refined to remove duplicates. The search process is regulated by the exploration-exploitation balance parameter ($\lambda$) and search iterations ($T$), tuned via grid search for efficiency and accuracy. We set $\lambda{=}0.5$ for prior-guided exploration before likelihood-driven exploitation and cap iterations at $T{=}3000$ for smaller datasets and $T{=}1000$ for larger ones to control computational cost. Re-ranking weights $\lambda_{a}$ and $\lambda_{o}$ are optimized within $[0.3,0.7]$ to balance action-object contributions. All experiments run on an NVIDIA RTX 3090 GPU, with an average inference time of 2 seconds per video.
@@ 6 Limitations, Discussion, and Future Work
While ProbRes significantly improves open-world egocentric activity recognition through structured search, challenges remain. Its reliance on Vision-Language Models (VLMs) inherits biases from pretraining data, occasionally leading to inefficient search trajectories. Structured priors help mitigate this, but the lack of semantic organization in text embeddings results in erratic jumps, suggesting that hyperbolic embeddings could enforce better locality. In higher openness levels (L2 and L3), LLM-generated search spaces enable scalable inference but require careful curation to avoid overly broad or noisy label distributions. Despite reducing VLM calls, further efficiency improvements—such as hierarchical search or adaptive refinement—are necessary for real-world deployment. ProbRes demonstrates that structured reasoning can replace brute-force enumeration in open-world settings. Integrating commonsense priors with probabilistic search lays the foundation for scalable and interpretable egocentric AI. Future work will explore additional cues, including gaze, human-object interactions, and scene context, to further enhance inference in unconstrained and dynamic real-world environments.
Acknowledgement. This research was supported in part by the US National Science Foundation grants IIS 2348689 and IIS 2348690. We thank Dr. Anuj Srivastava (FSU) and Dr. Sudeep Sarkar (USF) for their thoughtful feedback during the discussion about the project’s problem formulation and experimental analysis phase.

### v2

@@ Abstract
Open-world egocentric activity recognition poses a fundamental challenge due to its unconstrained nature, requiring models to infer unseen activities from an expansive, partially observed search space. We introduce ProbRes, a Probabilistic Residual search framework based on jump-diffusion that efficiently navigates this space by balancing prior-guided exploration with likelihood-driven exploitation. Our approach integrates structured commonsense priors to construct a semantically coherent search space, adaptively refines predictions using Vision-Language Models (VLMs) and employs a stochastic search mechanism to locate high-likelihood activity labels while minimizing exhaustive enumeration efficiently. We systematically evaluate ProbRes across multiple openness levels (L0–L3), demonstrating its adaptability to increasing search space complexity. In addition to achieving state-of-the-art performance on benchmark datasets (GTEA Gaze, GTEA Gaze+, EPIC-Kitchens, and Charades-Ego), we establish a clear taxonomy for open-world recognition, delineating the challenges and methodological advancements necessary for egocentric activity understanding. Our results highlight the importance of structured search strategies, paving the way for scalable and efficient open-world activity recognition. Code (in supplementary) will be shared publicly after review.
@@ 1 Introduction
Egocentric activity recognition in open-world settings presents a fundamental challenge in perception-driven reasoning, where an intelligent system must infer ongoing activities from a vast, unconstrained space of possibilities. Unlike closed-set classification, where predefined categories limit ambiguity, open-world recognition requires models to adaptively infer plausible activities, even when faced with unknown or unseen combinations of actions and objects. This challenge is particularly relevant for assistive robotics, wearable AI, and autonomous agents, where real-time understanding of human actions is crucial for interaction and decision-making. While multimodal foundation models such as Vision-Language Models (VLMs) have demonstrated remarkable zero-shot generalization, their reliance on exhaustive enumeration for inference makes them inefficient for large-scale open-world reasoning. To bridge this gap, structured knowledge is essential to move beyond naive brute-force enumeration, enabling scalable and efficient activity understanding in real-world applications.
@@
The increasing focus on open-world learning [26, 2, 13, 20] has led to ambiguity in defining openness, as different factors, objects, actions, and domains can independently contribute to it. Without a clear distinction, comparing methods and assessing generalization remain challenging. To address this, we define a hierarchy of openness levels that systematically capture increasing uncertainty in the search space (Figure 1). At L0, as in traditional zero-shot learning, all possible activity categories (action-object combinations) are predefined, but some may lack training examples. This setting benefits directly from Vision-Language Models (VLMs), which can generalize to unseen instances but operate within a bounded conceptual space. At L1, the atomic concept space — objects and actions — is known, but their composition (activities) is undefined, requiring models to reason about plausible pairings. At L2, the domain is known (e.g., cooking), but the underlying search space is unconstrained. Finally, at L3, no prior knowledge about the domain, actions, or objects is available, requiring models to construct the entire search space on the fly. As openness increases, exhaustive enumeration becomes intractable and requires structured priors for adaptive search.
@@
To address these, we introduce Probabilistic Residual Search (ProbRes), a structured search framework designed to navigate unconstrained activity spaces efficiently. Instead of relying solely on VLM-based likelihood estimation, ProbRes integrates structured priors from external knowledge sources to guide exploration, ensuring semantically meaningful search trajectories without supervision. As illustrated in Figure 2, ProbRes operates in three key phases: (1) exploration, where prior-driven sampling identifies plausible activity candidates; (2) exploitation, where the search refines predictions based on VLM likelihoods; and (3) residual refinement, which further optimizes results by re-ranking activities based on their decomposed action-object components. By balancing structured priors with data-driven reasoning, ProbRes efficiently locates high-likelihood activities while drastically reducing VLM queries.
@@
Our contributions are four-fold: (i) we propose ProbRes, a probabilistic residual search strategy that integrates structured priors with likelihood-driven inference, effectively balancing exploration and exploitation for open-world activity recognition, (ii) we introduce a clear hierarchical taxonomy (L0–L3) to systematically define and evaluate open-world recognition challenges, providing a principled foundation for future research, (iii) we show that ProbRes drastically reduces VLM queries while maintaining or improving recognition accuracy, demonstrating the importance of structured priors in large-scale search, and (iv) we highlight the limitations of current VLM-based text embeddings for search efficiency with qualitative analysis and highlight the need for improved semantic structuring.
@@ 2 Related Work
Egocentric video understanding has gained significant traction in the computer vision community in recent years due to its critical applications in areas such as imitation learning for robotic skills [45], virtual reality [21], and autonomous skill acquisition [22, 27, 4]. A wide range of tasks have been proposed to analyze egocentric video understanding, including procedure understanding, EgoPose [18, 19], proficiency estimation, question answering [15], gaze prediction [1, 29, 16], activity recognition [28], and summarization [33], driven by the development of large-scale datasets such as EgoExo4D [19], Ego4D [18], EPIC-Kitchens [10], Charades-Ego [46], GTEA Gaze [29], GTEA Gaze Plus [16], and Assembly101 [44].
@@
Despite these advancements, action recognition in egocentric videos remains challenging, particularly under conditions of rapid camera movement and significant occlusion, leading to issues with accuracy and generalization across different environments. While early approaches to egocentric action recognition were dominated by supervised learning methods, such as spatial-temporal modeling [49], motion-based representations [34], hand-object interactions [53, 62, 56], and time-series modeling [43], recent research has shifted towards self-supervised learning [59, 57], leveraging large-scale pretraining followed by fine-tuning for downstream tasks. Both paradigms remain data-driven and require extensive training to achieve optimal performance. Multi-modal LLM-based models such as MM-Ego [58] and GPT4Ego [9] have recently been explored for egocentric action recognition. Zero-shot learning [46, 60] offers a potential avenue to reduce dependence on labeled datasets, but its effectiveness is constrained by the inherent limitations of pre-trained models in reasoning about open-world activity compositions.
@@
Open-world understanding, particularly in egocentric activity recognition, remains an underexplored area. While significant progress has been made in open-world object detection, such as Open World DETR [13] and open-vocabulary object detection [20, 14], which leverage VLMs through prompting and knowledge distillation, challenges remain in handling novel class biases, as highlighted by UMB [55]. One of the first works addressing open-world egocentric activity recognition, KGL [2], employed a neuro-symbolic framework utilizing ConceptNet [47] for knowledge-driven inference. However, its reliance on object detectors for grounding concepts and the semantic gap between knowledge bases and video data limits its applicability. Chatterjee et al. [7] proposed a verb encoder and prompt-based object encoder for open-vocabulary action recognition in egocentric videos. More recently, ALGO [26] extended neuro-symbolic prompting to refine action-object affinities iteratively, combining object grounding with video foundation models. Other works explore related directions, such as incremental open-world action recognition [39] and skeleton-based methods for open-set action recognition [36].
@@
Vision-Language Models have gained prominence following the success of transformer-based LLMs such as BERT [12], RoBERTa [32], GPT series [40, 41, 5, 6], and open-source models like LLaMA [51, 52, 17] and DeepSeek [11]. Image-language models, including CLIP [42], DeCLIP [30], and ALIGN [23], leverage large-scale image-text pairs with contrastive learning [8, 24] to achieve strong zero-shot classification. Video-language models extend these capabilities to egocentric video understanding. EGO-VLP [31] employs separate video and text encoders, Hier-VL [3] captures multi-scale video information hierarchically and LAVILA [61] enhances video-text embeddings with LLM-generated narrations. EgoVLPv2 [38] integrates cross-modal fusion, while ALANAVLM [50] advances large-scale multimodal pretraining for embodied AI. However, these models are inherently constrained by the fixed search space defined during inference, limiting their generalization ability in open-world scenarios.
@@ 3 ProbRes: Jump Diffusion-based Reasoning
Overview. We aim to infer an activity label $a\in\mathcal{S}$ from a given video feature $v$. Unlike conventional closed-set recognition [49], where all activities are known during training, open-world settings require reasoning over a vast, partially observed search space, making exhaustive enumeration computationally prohibitive. Given a prior probability distribution $P_{\text{prior}}(a)$ capturing activity occurrence, a Vision-Language Model (VLM) likelihood estimator $P_{\text{likelihood}}(v{\mid}a)$, and a search space $\mathcal{S}$, the goal is to estimate the most probable activity efficiently:
@@
5:   $a_{t+1}\sim P_{\text{explore}}(a)$ $\triangleright$ Sample with exploration
[…]
@@
8:   $a_{t+1}\sim P_{\text{guided}}(a)$ $\triangleright$ Sample with exploitation
@@
[…]
@@
14:  $a\rightarrow(a_{\text{action}},a_{\text{object}})$ $\triangleright$ Decompose activity
15:  $S_{a}=v^{T}\phi(a_{\text{action}})$ $\triangleright$ Action Likelihood
16:  $S_{o}=v^{T}\phi(a_{\text{object}})$ $\triangleright$ Object Likelihood
@@ 3.1.2 Search Space Structuring
[…]
@@ 3.2.3 Concept Decomposition and Re-Ranking
Implementation Details. We use EGOVLP [31] and LAVILA [61] as VLM backbones for likelihood estimation, guided by structured priors from ConceptNet. For L2 and L3 evaluations, search spaces were generated using Gemini 2.0 Flash [37] and refined to remove duplicates. The search process is regulated by the exploration-exploitation balance parameter ($\lambda$) and search iterations ($T$), tuned via grid search for efficiency and accuracy. We set $\lambda{=}0.5$ for prior-guided exploration before likelihood-driven exploitation and cap iterations at T=3000T{=}3000 for smaller datasets and T=1000T{=}1000 for larger ones to control computational cost. Re-ranking weights $\lambda_{a}$ and $\lambda_{o}$ are optimized within $[0.3,0.7]$ to balance action-object contributions. All experiments run on an NVIDIA RTX 3090 GPU, with an average inference time of 2 seconds per video.
@@ 6 Limitations, Discussion, and Future Work
While ProbRes significantly improves open-world egocentric activity recognition through structured search, challenges remain. Its reliance on Vision-Language Models (VLMs) inherits biases from pretraining data, occasionally leading to inefficient search trajectories. Structured priors help mitigate this, but the lack of semantic organization in text embeddings results in erratic jumps, suggesting that hyperbolic embeddings could enforce better locality. In higher openness levels (L2 and L3), LLM-generated search spaces enable scalable inference but require careful curation to avoid overly broad or noisy label distributions. Despite reducing VLM calls, further efficiency improvements—such as hierarchical search or adaptive refinement—are necessary for real-world deployment. ProbRes demonstrates that structured reasoning can replace brute-force enumeration in open-world settings. Integrating commonsense priors with probabilistic search lays the foundation for scalable and interpretable egocentric AI. Future work will explore additional cues, including human-object interactions and scene context, to enhance inference in unconstrained.
Acknowledgement. This research was supported in part by the US National Science Foundation grants IIS 2348689 and IIS 2348690. We thank Dr. Anuj Srivastava (FSU), Dr. Sudeep Sarkar (USF), and the anonymous reviewers for their thoughtful feedback to help present this work better.